# Quest for optimal quantum resetting: protocols for a particle on a chain

### v3

@@ Abstract
In the classical context, it is well known that, sometimes, if the search does not find its target, it is better to start the process anew again, known as resetting. The quantum counterpart of resetting also indicates speeding up the detection process by eliminating the dark states, i.e., situations where the particle avoids detection. In this work, we introduce a most probable position resetting (MPR) protocol in which we reset the particle in a position where the probability of finding the particle could have been maximum, provided one would let the system evolve Unitarily in a given time window. In a tight-binding lattice model, there exists a 2-fold degeneracy (left and right) of the positions of maximum probability. The survival probability with optimal restart rate approaches zero (detection probability approaches one) when the particle is reset with equal probability on both sides. This protocol significantly reduces the optimal mean first-detected-passage time (FDT) and performs better even if the detector is far apart compared to the usual resetting protocols where the particle is brought back to the initial position. We propose a modified protocol, adaptive two-stage MPR, by making the associated probabilities of resetting to the right and left a function of resetting steps. In this protocol, we see a further reduction of the optimal mean FDT and improvement in the search process when the detector is far apart.
@@ I Introduction
Figure 1: Schematic of our proposed strategy based on the real data. We first place the particle at $j=0$, and the detector is at $\delta=10$, and let the system evolve till a time $t_{r}=3$ (resetting time) and then resets. The black dotted lines represent the probability distribution after a very short time of resetting. The resetting probability at the right (left) side peak is $p_{1}$ ($q_{1}$), which is denoted by a green (violet) dashed line. In the next steps, we follow the same strategy but with probability $p_{2}$, $p_{3}$, $\cdots$ to reset at the right peak and $q_{2}$, $q_{3}$, $\cdots$ to reset at the left peak, respectively. In the schematic, prob stands for probability. For the MPR protocol, we consider $p_{1}=p_{2}=p_{3}=\cdots=p$, while for the adaptive MPR protocol, we choose $p_{1}=p_{2}=\cdots=p_{R_{c}}=p_{I}$ and $p_{R_{c}+1}=p_{R_{c}+2}=\cdots=p_{F}$.
@@
In this work, our goal is to find out the optimal resetting schemes under repeated monitoring of a particle moving on a tight-binding lattice that can be used as an efficient protocol for the search process. In this context, it is essential to point out that there has been a plethora of work on the quantum walk on tight binding lattice without resetting Thiel et al. (2020b); Krovi and Brun (2006); Dhar et al. (2015a); Friedman et al. (2017a); Dubey et al. (2021); Thiel and Kessler (2020); Modak and Aravinda (2023); Thiel et al. (2018a). We propose a scheme according to which we do not reset the particle to its initial position; instead, the particle is reset to sites where the probability of finding the particle would have been maximum, provided the evolution of the quantum state has not been interrupted by repetitive measurements, called most probable position resetting (MPR). Given the tight-binding lattice and our choice of initial state, such sites are doubly degenerate. We associate probabilities of resetting to each of those two sites and reset accordingly (see Fig. 1 for the schematic of our protocols). We like to emphasize that in the case of the unbiased classical random walk, the probability distribution will always have a peak in the initial position; hence, the quantum counterpart of the protocol is the same as the initial position resetting (IPR) Yin and Barkai (2023b). However, there is no classical counterpart to the MPR protocol. We find that MPR protocol reduces the mean FDT significantly in comparison to the previously studied IPR protocols for optimal resetting time. Moreover, we introduce a two-stage adaptive most probable position resetting (Adaptive two-stage MPR) protocol by carefully varying the probability of choosing one of the double degenerate most likely visited sites in each resetting step. Further reduction in mean FDT is observed in this protocol. The optimal mean FDT increases linearly with the increasing distance of the detector from the center. The slope of this linear dependence is smaller for MPR and adaptive MPR than the previously studied resetting protocols Yin and Barkai (2023b), which signifies the advantage of MPR and adaptive two-stage MPR over IPR when the detector is far apart from the center.
@@ III.2 Most Probable Position Resetting(MPR) protocol $\mathcal{S}^{p}_{x_{M}}$
In this section, we propose a resetting protocol where, instead of resetting to the initial position, i.e., at the site $x_{0}$ every time, we reset the particle to the site where the probability of finding it could have been maximum in the absence of measurement. We start from an initial state $|x_{0}\rangle$. If the wave function is allowed to evolve Unitarily under the Hamiltonian $H$ for time $t_{r}=r\tau$, the probability of finding the particle at any site $j$ at time $t_{r}$ is $P(j,x_{0},r\tau)$ (see Eq. (2)). In that case, the site where the particle most likely will be after time $t_{r}$ corresponds to the value of $j$ for which $P(j,x_{0},r\tau)$ is maximum. We identify that site as $j^{*}_{1}$, and can write,
@@
Now, we reset the particle at $j_{1}^{*}$ (instead of $x_{0}$). Similarly, in the next steps, one resets the system at,
@@
For the tight-binding Hamiltonian $H$ (where all the sites are identical), it turns out that if the resetting time $t_{r}$ is larger than some critical value (see Appendix. F), the peak position of the traveling wave front $j^{*}_{k}$ will always be doubly degenerate, while below the critical $t_{r}$, $j^{*}_{k}=x_{0}$ $\forall k\in[1,R]$. For example, for a given $t_{r}$, if
@@
The above equation represents, at the $R$’th resetting step we have possible resetting positions at [$x_{0}+(R-2m)\Delta$](occurring $\binom{R}{m}$ times), where $m=0,1,\cdots R$, the set of all possible resetting positions at $R$-th resetting step, we are referring as $j_{R}^{*}$, and here $\Delta=|j^{*}_{1}-x_{0}|\geq 0$.
Note that for $\Delta=0$, which corresponds to very small $t_{r}$, the protocol is the same as the IPR protocol $\mathcal{S}_{x_{0}}$. However, for non-zero $\Delta$s, $j^{*}_{k+1}=j^{*}_{k}\pm\Delta$. We associate a real number $p$ with our scheme such that $0\leq p\leq 1$, and $p$ stands for the probability of resetting at the site $j^{*}_{k}+\Delta$ (and $q=1-p$ is the probability of resetting at $j^{*}_{k}-\Delta$) $\forall k\in[1,R]$. It means if we initialize the particle at $x_{0}$, then in the 1st step we reset the particle at $x_{0}+\Delta$ with probability $p$ and at $x_{0}-\Delta$ with probability $q$, further in the 2nd step we reset the particle at $x_{0}+2\Delta$ with probability $p^{2}$, at $x_{0}-2\Delta$ with probability $q^{2}$, and at $x_{0}$ with probability $2pq$, and so on (see Fig. 1). Important to emphasize that in our protocol, we are not keeping track of the resetting history of the particle in the previous steps, e.g., in the 2nd step of our protocol, in order to reset the particle at $x_{0}+2\Delta$ ($x_{0}-2\Delta$); it is not necessary that previously the particle was reset at site $x_{0}+\Delta$ ($x_{0}-\Delta$). Independent of where the particle was reset in the 1st step, in the 2nd step, we reset them at $x_{0}+2\Delta$ with probability $p^{2}$, at $x_{0}-2\Delta$ with probability $q^{2}$, and at $x_{0}$ with probability $2pq$. We have also studied a scenario where we track the previous resetting trajectories of the particle (which we refer to as path-dependent MPR protocol); results remain qualitatively the same as what we find here, which are described in Appendix. E. However, in our case, given that quantum dynamics are affected by repetitive measurement, the dynamics do not remain unitary. As shown in Appendix. C, depending on the detector’s location and resetting time, the wave packet’s peak position can be quite different from the wave packet of measurement-free evolution. That can potentially make the path-dependent protocols harder to implement than the path-independent ones, where the knowledge of resetting history is not required.
If in the whole process, we restart the system $R$ times in the positions $j_{1}^{*}$, $j_{2}^{*}$,……$j_{R}^{*}$ respectively, where $j_{R}^{*}$ represents the set of all possible resetting positions at $R$-th resetting step, then one can modify Eq. (11) according to MPR protocol $S^{p}_{x_{M}}$ and can write,
@@
$\tilde{F}_{{n}}^{R}$ represents the first detection probability in the $n-th$ measurement, summing over the contributions of starting from all possible sites in the $R-th$ resetting step with the corresponding probability of being that particular site. Inserting the modified expression of $F_{n}(r)$ into Eq.(12) and Eq. (14), one can obtain the total probability of detection $P^{r}_{det}$ up to time $n\tau$ and the mean FDT ${\langle t_{f}\rangle}_{r}$ for the protocol $\mathcal{S}^{p}_{x_{M}}$. The results of the above expressions can be interpreted as, at a resetting step $R$, we reset at any of the positions from set $j_{R}^{*}$ with the appropriate probability. Again in the next step, we reset at any of the possible positions from the set $j_{R+1}^{*}$ with the corresponding probability without tracking the previous reset trajectories. We verify the correctness of the above equations by performing a numerical experiment (see Appendix D for details).
@@
We introduce another protocol, which is a modified version of MPR protocol $\mathcal{S}^{p}_{x_{M}}$, where we had the probability $p$ of resetting at the site $j^{*}_{k}+\Delta$ remains the same in each resetting steps. On the other hand, here we propose a scheme where the probability $p$ can also vary in each resetting step, e.g. one can think of a situation such that $p_{1}$ ($q_{1}$) is the probability of resetting at the site $j^{*}_{1}=x_{0}+\Delta$ ($j^{*}_{1}=x_{0}-\Delta$), $p_{2}$ ($q_{2}$) is the probability of resetting at the site $j^{*}_{1}+\Delta$ ($j^{*}_{1}-\Delta$), and so on. In this case, one can write,
@@
Given the numbers of individual terms increase exponentially with $R$ in Eq. (19), here we consider a simpler scheme, i.e. we associate only two real numbers $p_{I}$ ($q_{I}=1-p_{I}$) and $p_{F}$ ($q_{F}=1-p_{F}$) with our scheme, and they stand for the probability of resetting at the site $j^{*}_{k}+\Delta$ ( $j^{*}_{k}-\Delta)\quad\forall k\leq R_{c}$ and $\forall k>R_{c}$ respectively. We refer to this scheme as a two-stage Adaptive MPR protocol $\mathcal{S}^{p_{I},p_{F}}_{x_{M}}$. One is free to choose the optimal value for $p_{I}$, $p_{F}$, and $R_{c}$ to speed up the detection process. We choose $p_{F}=0.5$ (the reason is discussed in the next section), and $p_{I}$ can be thought of as an adaptive /confidence parameter. Suppose, based on some previous knowledge, if one thinks that more likely the search object is on the right side of the lattice, $p_{I}$ can be chosen $>0.5$. Based on the confidence level, one will choose $p_{I}$ accordingly.

### v4

@@ Abstract
In the classical context, it is well known that, sometimes, if the search does not find its target, it is better to start the process anew again, known as resetting. The quantum counterpart of resetting also indicates speeding up the detection process by eliminating the dark states, i.e., situations where the particle avoids detection. In this work, we introduce the most probable position resetting (MPR) protocol in which, at a given resetting step, resets are done with certain probabilities to the set of possible peak positions (where the probability of finding the particle is maximum) that could occur because of the previous resets and followed by uninterrupted unitary evolution, irrespective of which path was taken by the particle in previous steps. In a tight-binding lattice model, there exists a 2-fold degeneracy (left and right) of the positions of maximum probability. The survival probability with optimal restart rate approaches zero (detection probability approaches one) when the particle is reset with equal probability on both sides path independently. This protocol significantly reduces the optimal mean first-detected-passage time (FDT) and performs better even if the detector is far apart compared to the usual resetting protocols where the particle is brought back to the initial position. We propose a modified protocol, an adaptive two-stage MPR, by making the associated probabilities of going to the right and left a function of steps. In this protocol, we see a further reduction of the optimal mean FDT and improvement in the search process when the detector is far apart.
@@ I Introduction
Figure 1: Schematic of our proposed strategy based on the real data. First, we initialize the particle at $j=0$ site and let the wave function evolve for the time $t_{r}=3$ without any measurement interruptions. We see two peaks at $j=5$ and $j=-5$. Hence, we reset at positions $j=+5$ or $j=-5$. Now, as we have degeneracy in the reset positions, we associate probabilities $p_{1}$ and $q_{1}$ for reaching those right and left resetting positions by random walk starting from $j=0$. So, for the first resetting step, we reset either at $j=+5$ with probability $p_{1}$ or at $j=-5$ with probability $q_{1}$ (upper panel). Now, when we reset for the $2$nd time, we do not carry the history of the $1$st reset. Starting from $j=+5$, the next peaks are at $j=0,10$, and if it starts at $j=-5$, the next peaks are at $j=0,-10$. We have a total of three possible resetting positions at $j=0,+10,-10$ (middle panel). We associate the probabilities of choosing those peaks as going to those peak positions starting from $j=0$ by two steps doing a random walk, where the $1$st random walk step has the probability of going left and right as $q_{1}$ and $p_{1}$, respectively, and the $2$nd step has the probability $q_{2}$ and $p_{2}$ respectively. Hence, these three peaks at $j=\{-10,0,+10$} are having the resetting probabilities $\{q^{2},2pq,p^{2}\}$. Similarly, in the $3rd$ step, we have $4$ peaks at $j=-15,-5,5,15$. We don’t carry the memory of the previous steps and simply associate the probability of choosing the peaks as going to those peak positions from $j=0$ by random walk using three steps, where the third step of the random walk has the probability of going left and right as $q_{3}$ and $p_{3}$ (lower panel). We consider $p_{1}=p_{2}=p_{3}=\cdots=p$ for the MPR protocol, while for the adaptive MPR protocol, we choose $p_{1}=p_{2}=\cdots=p_{R_{c}}=p_{I}$ and $p_{R_{c}+1}=p_{R_{c}+2}=\cdots=p_{F}$.
@@
In this work, our goal is to find out the optimal resetting schemes under repeated monitoring of a particle moving on a tight-binding lattice that can be used as an efficient protocol for the search process. In this context, it is essential to point out that there has been a plethora of work on the quantum walk on tight binding lattice without resetting Thiel et al. (2020b); Krovi and Brun (2006); Dhar et al. (2015a); Friedman et al. (2017a); Dubey et al. (2021); Thiel and Kessler (2020); Modak and Aravinda (2023); Thiel et al. (2018a). We propose a scheme according to which we do not reset the particle to its initial position; instead, at a given resetting step, resets are done with certain probabilities to the set of possible peak positions (where the probability of finding the particle is maximum) that could occur because of the previous resets and followed by uninterrupted unitary evolution, irrespective of which path was taken by the particle in previous steps, called the most probable position resetting (MPR). As an example, we demonstrate our protocols in the form of a schematic based on the real data in Fig. 1. We like to emphasize that in the case of the unbiased classical random walk, the probability distribution will always have a peak in the initial position; hence, the quantum counterpart of the protocol is the same as the initial position resetting (IPR) Yin and Barkai (2023b). However, there is no classical counterpart to the MPR protocol. We find that the MPR protocol reduces the mean FDT significantly in comparison to the previously studied IPR protocols for optimal resetting time. Moreover, we introduce a two-stage adaptive most probable position resetting (Adaptive two-stage MPR) protocol by carefully varying the probability of going left and right in each step. Further reduction in mean FDT is observed in this protocol. The optimal mean FDT increases linearly with the increasing distance of the detector from the center. The slope of this linear dependence is smaller for MPR and adaptive MPR than the previously studied resetting protocols Yin and Barkai (2023b), which signifies the advantage of MPR and adaptive two-stage MPR over IPR when the detector is far apart from the center.
@@ III.2 Most Probable Position Resetting(MPR) protocol $\mathcal{S}^{p}_{x_{M}}$
In this section, we propose a resetting protocol where, instead of resetting to the initial position, i.e., at the site $x_{0}$ every time, in each resetting step, resets are done with certain probabilities to the set of possible peak positions that could occur because of the previous resets and followed by uninterrupted unitary evolution, irrespective of which path was taken by the particle in previous steps. We start from an initial state $|x_{0}\rangle$. If the wave function is allowed to evolve unitarily under the Hamiltonian $H$ for time $t_{r}=r\tau$, the probability of finding the particle at any site $j$ at time $t_{r}$ is $P(j,x_{0},r\tau)$ (see Eq. (2)). In that case, the site where the particle most likely will be after time $t_{r}$ corresponds to the value of $j$ for which $P(j,x_{0},r\tau)$ is maximum. We identify that site as $j^{*}_{1}$, and can write,
@@
In the next steps, one has,
@@
For the tight-binding Hamiltonian $H$ (where all the sites are identical), it turns out that if the resetting time $t_{r}$ is larger than some critical value (see Appendix. F), the peak position of the traveling wave front will always be doubly degenerate, while below the critical $t_{r}$, $j^{*}_{k}=x_{0}$ $\forall k\in[1,R]$. For example, for a given $t_{r}$, if
@@
The above equation represents, at the $R$’th resetting step we have possible resetting positions at [$x_{0}+(R-2m)\Delta$](occurring $\binom{R}{m}$ times), where $m=0,1,\cdots R$, the set of all possible resetting positions at $R$-th resetting step, we are referring as $j_{R}^{*}$, and here $\Delta=|j^{*}_{1}-x_{0}|\geq 0$. Note that for $\Delta=0$, which corresponds to very small $t_{r}$, the protocol is the same as the IPR protocol $S_{x_{0}}$. At a given resetting step, we reset at any of the possible resetting positions from the set $j_{R}^{*}$ irrespective of the fact that where it was reset in the previous steps, that makes our protocol path-independent. Also, we associate probabilities for choosing a reset position at a given resetting step $R$ from the set of available resetting position $j_{R}^{*}$ as the probability of going to that particular position doing random walk by $R$ step starting from $x_{0}$. For the MPR protocol, for every random walk step, we associate probabilities $p$ and $q$ for right and left, respectively. It means if we initialize the particle at $x_{0}$, then in the 1st step we reset the particle at $x_{0}+\Delta$ with probability $p$ and at $x_{0}-\Delta$ with probability $q$, further in the 2nd step we reset the particle at $x_{0}+2\Delta$ with probability $p^{2}$, at $x_{0}-2\Delta$ with probability $q^{2}$, and at $x_{0}$ with probability $2pq$, and so on (see Fig. 1). Important to emphasize that in our protocol, we are not keeping track of the resetting history of the particle in the previous steps, e.g., in the 2nd step of our protocol, in order to reset the particle at $x_{0}+2\Delta$ ($x_{0}-2\Delta$); it is not necessary that previously the particle was reset at site $x_{0}+\Delta$ ($x_{0}-\Delta$). Independent of where the particle was reset in the 1st step, in the 2nd step, we reset them at $x_{0}+2\Delta$ with probability $p^{2}$, at $x_{0}-2\Delta$ with probability $q^{2}$, and at $x_{0}$ with probability $2pq$. We have also studied a scenario where we track the previous resetting trajectories of the particle (which we refer to as path-dependent MPR protocol); results remain qualitatively the same as what we find here, which are described in Appendix. E. However, note that the path-dependent MPR protocol is difficult to analyze theoretically as compared to the path-independent one. Hence, we have restricted ourselves to the path-independent protocols in the main text.
If in the whole process, we restart the system $R$ times in the positions taken from the sets $j_{1}^{*}$, $j_{2}^{*}$,……$j_{R}^{*}$ respectively, where $j_{R}^{*}$ represents the set of all possible resetting positions at $R$-th resetting step, then one can modify Eq. (11) according to MPR protocol $S^{p}_{x_{M}}$ and can write,
@@
$\tilde{F}_{{n}}^{R}$ represents the first detection probability in the $n-th$ measurement, summing over the contributions of starting from all possible sites in the $R-th$ resetting step with the corresponding probability of being that particular site.
Let us validate the above expressions Eq. (17) and Eq. (18) with an example of $n=3,r=1,R=2$. For the $1st$ resetting step, possible peaks are at $x_{0}+\Delta$ and $x_{0}-\Delta$, and associated resetting probabilities are $p$ and $q$, respectively. In the $2nd$ step, the possible peaks are at $x_{0}+2\Delta,x_{0},x_{0}-2\Delta$ with resetting probabilities $p^{2},2pq,q^{2}$ respectively independent of where the particle was reset in the 1st step. Hence, in the expression of $F_{3}(1)$, there will be contributions from $6$ possible paths: 1) resetted at $x_{0}+\Delta$ in the 1st step, and then in the 2nd step at $x_{0}+2\Delta$, 2) resetted at $x_{0}+\Delta$ in the 1st step, and then in the 2nd step at $x_{0}$, 3) resetted at $x_{0}+\Delta$ in the 1st step, and then in the 2nd step at $x_{0}-2\Delta$. 4) resetted at $x_{0}-\Delta$ in the 1st step, and then in the 2nd step at $x_{0}+2\Delta$, 5) resetted at $x_{0}-\Delta$ in the 1st step, and then in the 2nd step at $x_{0}$, and 6) resetted at $x_{0}-\Delta$ in the 1st step, and then in the 2nd step at $x_{0}-2\Delta$. Mathematically, it reads as,
$F_{3}(1)=$ $(1-|\phi_{1}^{x_{0}}|^{2})p(1-|\phi_{1}^{x_{0}+\Delta}|^{2})p^{2}(|\phi_{1}^{x_{0}+2\Delta}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})p(1-|\phi_{1}^{x_{0}+\Delta}|^{2})2pq(|\phi_{1}^{x_{0}}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})p(1-|\phi_{1}^{x_{0}+\Delta}|^{2})q^{2}(|\phi_{1}^{x_{0}-2\Delta}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})q(1-|\phi_{1}^{x_{0}-\Delta}|^{2})p^{2}(|\phi_{1}^{x_{0}+2\Delta}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})q(1-|\phi_{1}^{x_{0}-\Delta}|^{2})2pq(|\phi_{1}^{x_{0}}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})q(1-|\phi_{1}^{x_{0}-\Delta}|^{2})q^{2}(|\phi_{1}^{x_{0}-2\Delta}|^{2})$
$=$ $(1-|\phi_{1}^{x_{0}}|^{2})(1-p|\phi_{1}^{x_{0}+\Delta}|^{2}-q|\phi_{1}^{x_{0}-\Delta}|^{2})$
$~{}~{}~{}\cdot\big{[}p^{2}(|\phi_{1}^{x_{0}+2\Delta}|^{2})+2pq(|\phi_{1}^{x_{0}}|^{2})+q^{2}(|\phi_{1}^{x_{0}-2\Delta}|^{2})\big{]}$
$=$ $(1-\tilde{P}^{0}_{det}(1))(1-\tilde{P}^{1}_{det}(1))\tilde{F}_{1}^{2}$
This is identical to what one obtained from the Eq. (17). See Appendix.D for the numerical validation in case of large $R$. On the other hand, for the path-dependent protocol, one would have contributions from only 4 paths: 1) resetted at $x_{0}+\Delta$ in the 1st step, and then in the 2nd step at $x_{0}+2\Delta$, 2) resetted at $x_{0}+\Delta$ in the 1st step, and then in the 2nd step at $x_{0}$, 3) resetted at $x_{0}-\Delta$ in the 1st step, and then in the 2nd step at $x_{0}$, and 4) resetted at $x_{0}-\Delta$ in the 1st step, and then in the 2nd step at $x_{0}-2\Delta$. Hence, mathematically, $F_{3}(1)$ for path-dependent protocol will read as,
$F_{3}(1)=$ $(1-|\phi_{1}^{x_{0}}|^{2})p(1-|\phi_{1}^{x_{0}+\Delta}|^{2})p(|\phi_{1}^{x_{0}+2\Delta}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})p(1-|\phi_{1}^{x_{0}+\Delta}|^{2})q(|\phi_{1}^{x_{0}}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})q(1-|\phi_{1}^{x_{0}-\Delta}|^{2})p(|\phi_{1}^{x_{0}}|^{2})$
$+(1-|\phi_{1}^{x_{0}}|^{2})q(1-|\phi_{1}^{x_{0}-\Delta}|^{2})q(|\phi_{1}^{x_{0}-2\Delta}|^{2}).$
Note that this expression is different from the results obtained for the path-independent protocol using Eq. (17)(see Appendix. E for more details). Unlike path-independent protocol, it is not possible to get a nice expression like Eq.(17) in the case of path-dependent protocol. That makes the path-independent protocol easier to analyze theoretically as compared to the path-dependent one.
Inserting the modified expression of $F_{n}(r)$ into Eq.(12) and Eq. (14), one can obtain the total probability of detection $P^{r}_{det}$ up to time $n\tau$ and the mean FDT ${\langle t_{f}\rangle}_{r}$ for the protocol $\mathcal{S}^{p}_{x_{M}}$. The results of the above expressions can be interpreted as, at a resetting step $R$, we reset at any of the positions from set $j_{R}^{*}$ with the appropriate probability. Again in the next step, we reset at any of the possible positions from the set $j_{R+1}^{*}$ with the corresponding probability without tracking the previous reset trajectories. We verify the correctness of the above equations by performing a numerical experiment (see Appendix D for details).
@@
Finally, we summarize our protocols in $4$ simple steps,
Step 1: We initialize the particle at the center of the lattice at $x_{0}$ and keep the detector at site $\delta$, which clicks in the interval $\tau,2\tau,3\tau,\cdots,n\tau,\cdots$.
Step 2: To compute the first detection probability $F_{n}(r)$ in the $n$-th measurement, first we fix a resetting time $r\tau$, which automatically fixes the maximum number of resetting steps are allowed, i.e, $R=int[\frac{n-1}{r}]$. Identify the distance $\Delta$ of the peaks in the time-evolved probability density from the initial position of the particle at time $r\tau$.
Step 3: Keep resetting the particles at time intervals of $r\tau,2r\tau,\cdots,Rr\tau$, at the following sites $\{x_{0}+\Delta,x_{0}-\Delta\},\{x_{0}+2\Delta,x_{0},x_{0}-2\Delta\},\cdots,\{[x_{0}+(R-2m)\Delta]\},\cdots$, where $0\leq m\leq R,m\in\mathbb{Z}$ respectively with probabilities $\{p,q\},\{p^{2},2pq,q^{2}\},\cdots,\{\binom{R}{m}p^{R-m}(1-p)^{m}\}$.
Step 4: Calculate the first detection probability $F_{n}(r)$ in the $n$-th measurement, taking the average over all possible realizations, mathematically, which is represented by Eq.(17).
@@
We introduce another protocol, which is a modified version of MPR protocol $\mathcal{S}^{p}_{x_{M}}$, where we had the probability $p$ ($q$) of going at right (left) remains the same in each step. On the other hand, here we propose a scheme where the probability $p$ can also vary in each random walk step in a given resetting step, e.g. one can think of a situation where in the $1st$ resetting step, we have the probability of resetting at $x_{0}+\Delta$ and $x_{0}-\Delta$ as $p_{1}$ and $q_{1}$ respectively. In the $2nd$ resetting step, the possible positions are $x_{0}+2\Delta,x_{0},x_{0}-2\Delta$. In this process, starting from $x_{0}$ (center of the lattice), we consider the probability of going to the left and right in the first random walk step is $p_{1}$ and $q_{1}$, and in the second random walk step, these probabilities become $p_{2}$ and $q_{2}$. The probability of reaching at $x_{0}+2\Delta,x_{0},x_{0}-2\Delta$ are $p_{1}p_{2},p_{1}q_{2}+q_{1}p_{2},q_{1}q_{2}$ respectively. A similar rule is applicable for all the resetting steps (see fig. 1). In this case, one can write,
@@
Given the numbers of individual terms increase exponentially with $R$ in Eq. (19), here we consider a simpler scheme, i.e. we associate only two real numbers $p_{I}$ ($q_{I}=1-p_{I}$) and $p_{F}$ ($q_{F}=1-p_{F}$) with our scheme. Previously, for the MPR protocol, the probability of resetting to a possible peak at the $R-th$ resetting step was taken as the probability of reaching that peak in $R$ steps starting from the site $x_{0}$ doing a random walk. According to the adaptive two-stage MPR protocol, we consider that among those $R$ random walk steps, $R_{c}$ number of steps have probability $p_{I}$($q_{I}$) of going right(left), and $R-R_{c}$ steps have probability $p_{F}$($q_{F}$) of going right(left). One is free to choose the optimal value for $p_{I}$, $p_{F}$, and $R_{c}$ to speed up the detection process. We choose $p_{F}=0.5$ (the reason is discussed in the next section), and $p_{I}$ can be thought of as an adaptive /confidence parameter. Suppose, based on some previous knowledge, if one thinks that more likely the search object is on the right side of the lattice, $p_{I}$ can be chosen $>0.5$. Based on the confidence level, one will choose $p_{I}$ accordingly.